OpenStudy (anonymous):

find the derivative of this equation (2x-1)^2 (3x+5)^4

OpenStudy (anonymous):

usee dat product rule wit dat chain rule

OpenStudy (anonymous):

or expand your choice

OpenStudy (agent0smith):

derivative of \[f(x)g(x) = f'(x)g(x) + f(x)g'(x)\]

OpenStudy (anonymous):

\[D[(u)^n]=n(u)^{n-1}u'\]]

OpenStudy (anonymous):

or \[D[f(g(x))=f'(x)g'(x)\]

OpenStudy (agent0smith):

let f(x) = (2x-1)^2 then f'(x) = 2(2x-1)*2 Then let g(x) = (3+5)^4 (i'm assuming an x should be in there) now find g'(x)